[Hallc_running] [New Logentry] RC Update Sat Feb 29

brads at jlab.org brads at jlab.org
Sat Feb 29 16:55:01 EST 2020

Logentry Text:
<h2><span class="mw-headline"><strong><a class="new" title="RC Meeting and Run Plan, Saturday, February 29th 2020" href="https://hallcweb.jlab.org/wiki/index.php?title=RC_Meeting_and_Run_Plan,_Saturday,_February_29th_2020">RC Meeting and Run Plan, Saturday, February 29th 2020</a></strong></span></h2><h2><span class="mw-headline"><strong>There will be no RC meeting Sat or Sun.</strong></span></h2><h1><span id="General_updates" class="mw-headline">General updates</span></h1><ul><li>Poor beam late Friday swing traced to 1L11, EES folks came in and were able to recover some cavities<ul><li>Subsequent beam was better</li></ul></li><li>Switched to Transverse polarization early-swing on Friday: seems to be good!<ul><li><a class="external free" href="https://logbooks.jlab.org/entry/3794040" rel="nofollow">https://logbooks.jlab.org/entry/3794040</a>; Note that the calib. constant for transverse needs to be remeasured</li><li>1st meas: 49.8% <a class="external autonumber" href="https:!
 //logbooks.jlab.org/entry/3794045" rel="nofollow">[1]</a>, 2nd meas: 49.2% <a class="external autonumber" href="https://logbooks.jlab.org/entry/3794079" rel="nofollow">[2]</a>, last meas. failed <a class="external autonumber" href="https://logbooks.jlab.org/entry/3794242" rel="nofollow">[3]</a></li><li>Rotated back to longitudinal early Sat. OWL (pol. still good) and continued with long. production data taking</li></ul></li><li>Plot of NMR data taken on Big Brother: <a class="external free" href="https://logbooks.jlab.org/entry/3794600" rel="nofollow">https://logbooks.jlab.org/entry/3794600</a> (prelim. data only)</li><li>Slight Beam energy shift when Hall A went offline (lock moved to 2A?) <a class="external autonumber" href="https://logbooks.jlab.org/entry/3794454" rel="nofollow">[4]</a></li><li>USB-ethernet bridge that was replaced on Friday seemed to fail Sat. morning <a class="external autonumber" href="https://logbooks.jlab.org/entry/3794496" rel="nofollow">[5]</a><ul!
 ><li>Took escorted access and swapped back (the new bridge did!
  not help with the initial disconnect problem anyway)</li></ul></li></ul><ul><li>Generally good production running during OWL into Sat day/swing.</li></ul><h2><span id="Notes_for_the_Future" class="mw-headline">Notes for the Future</span></h2><ul><li>Beam studies to understand what is going on with the machine are planned for Tuesday DAY shift<ul><li>We will try to get in in the morning (RC will talk to PD and RadCon -- may need to open target enclosure)</li><li>Details being worked out, but it'll be a full 7--8 hour shift</li><li>EPR and pNMR measurements/work are planned during this downtime</li><li>There are portions of the Beam Studies program that would involve beam delivery into Hall C (no target), so limited opportunity for Access only</li><li>If possible, we would like to do our access in the morning</li><li>May need to get into the Target enclosure (RadCon support)</li></ul></li><li>Injector spot move has been pushed off for at least another week (2--3 hour down)<u!
 l><li>Not much beam over the last several days...</li><li>Hope is to push it off to the Hall C down that starts March 13, and then just do a reactivation that would carry us through May 6<ul><li>Not sure how it will work out, but goal will be to minimized downs and maximize integrated current to Hall(s)</li></ul></li></ul></li></ul><ul><li>'Filler' runs for next spin up / target-down period:<ul><li>1 hour run with target out for measure of the background just from the beryllium windows</li></ul></li></ul><h2><span id="Pending_Issues" class="mw-headline">Pending Issues</span></h2><ul><li>Transverse NMR measurement that failed <a class="external autonumber" href="https://logbooks.jlab.org/entry/3794242" rel="nofollow">[6]</a></li><li>pNMR issues</li><li>EPR problems <a class="external autonumber" href="https://logbooks.jlab.org/entry/3793821" rel="nofollow">[7]</a><ul><li>Going forward for EPR, I'd try (1) checking the continuity of the cable leading from the RF amplifier to !
 the EPR coil to see if it's stable, (2) measure the current going throu!
 gh the EPR coil, (3) cleaning the single D2 filter (it was very dirty) and confirming the APD gain is indeed set to maximum, measure the DC light output (currently a measly 3.6 mV), then trying another FM sweep. If that doesn't improve it, then (4) remove the D2 filter, measure the DC light output, perform another FM scan.</li></ul></li><li>Kepco power supply commissioning/installation (Todd, Bill)</li><li>Update Transverse NMR calibration coeff.</li><li>Persistent BLM & ion chamber trips<ul><li>General activation, weird BLM issues (<a class="external free" href="https://logbooks.jlab.org/entry/3792326" rel="nofollow">https://logbooks.jlab.org/entry/3792326</a>), Ad-hoc ion chamber threshold adjustments (<a class="external free" href="https://logbooks.jlab.org/entry/3792095" rel="nofollow">https://logbooks.jlab.org/entry/3792095</a>)</li><li><strong>NOTE</strong>: Shift crew should <em>NOT</em> give authorization to MCC change ion chamber thresholds. If they ask, tell O!
 perator to contact an appropriate expert. Make a log entry if trips persist, or call the RC if it becomes severe (for example, >20 trips/hour is getting intolerable)</li></ul></li></ul><h2><span id="Opportunistic_Access_Jobs" class="mw-headline">Opportunistic Access Jobs</span></h2><ul><li>Target work during Beam Studies (see above)<ul><li>pNMR troubleshooting</li><li>EPR issues <a class="external autonumber" href="https://logbooks.jlab.org/entry/3793821" rel="nofollow">[8]</a></li><li>Transverse NMR?<a class="external autonumber" href="https://logbooks.jlab.org/entry/3794242" rel="nofollow">[9]</a></li></ul></li></ul><h2><span id="Target_coil_rotation_planned_for_March_13--20" class="mw-headline"><strong>Target coil rotation planned for March 13--20</strong></span></h2><ul><li><ul><li>Nominal 1-week downtime in Hall C to rotate target coils, install new cell, and transition to d2n program</li><li>Status Update for target cell <em>Austin</em>:<ul><li>Wall and window thic!
 knesses done</li><li>Only 'Reduced set' of density measurements availab!
 le now, will continue measurements this week (issues with laser at UVa?)</li><li>It would be best for it to arrive at JLab on March 10 (at the latest) to allow locals to mount cell in oven plate before they get involved with target work in Hall.</li></ul></li><li>Lots of preplanning to be completed (RadCon, Techs, S&A, Target group, etc)</li><li>RW-II training for JP, Arun, Brad (Complete the training documentation and take the RW-II test if you haven't already done so)</li></ul></li></ul>


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3794662
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200229/6f1d8406/attachment.html>

More information about the Hallc_running mailing list